anonomys:

The graph of f(x) = 2x + 1 is shown below. Explain how to find the average rate of change between x = 0 and x = 2.

ThisGirlPretty:

@Nnesha

ThisGirlPretty:

@Ultrilliam

Ultrilliam:

Hold up one sec

Ultrilliam:

Welcome to QC BTW!

Nnesha:

f(x) can be written as y \[\huge\rm y=2^x+1\] first of all we need to find y values that corresponds to x=0,2 substitute x for 0 and 2 to find y another word for *average rate of change* is *slope * formula to find slope is \[\large\rm \frac{ y_2-y_1 }{ x_2-x_1}\]

RektbyFaith:

Hey there buddy =) Welcome to QC :) So what we need to do is first use our Formula from when we were in geometry Which is \[Y_2-Y_1/X_2-X_1\] This formula is the formula for the average rate of change for two points. Now what we need to do is use a new formula, Which is now \[F_b - F_a /b-a\] Try to use this formula to solve your answer.

Nnesha:

\[y=2^x+1\]\[\large\rm \color{Red}{y}=2^\color{blue}{0}+1=1+1\] (anything raised to the zero power is equal to 1 when x =0 y =2 so when x=2 y=what?

Nnesha:

how did you get that ?

Nnesha:

y=2 when x=0 now replace x with 2 to find the y value when x=2

Nnesha:

\[y=2^\color{blue}{x}+1\]\[\large\rm \color{Red}{y}=2^\color{blue}{0}+1\]\[\large\rm \color{Red}{y}=1+1\]\[\large\rm \color{Red}{y}=2\] repeat the steps. This time replace `x` with `2`

Nnesha:

yes correct! so when x = 0 ,y=2 and when x =2 , y =5 we can write it as order pair `(0,2),(2,5)` now use the slope formula for the slope \[\large\rm (\color{blue}{x_1} ,\color{red}{y_1})(\color{blue}{x_2},\color{Red}{y_2})\] \[\large\rm slope=\frac{\color{Red}{ y_2 - y_1} }{\color{blue}{ x_2-x_1} }\]

Nnesha:

correct. good job

Nnesha:

np

Nnesha:

aww thanks for the nice words. appreciate 't!! o^_^o :=))